OpenStudy (kira_yamato):

Gas A exerts a pressure of 1.00 atm in a 1.00dm^3 container. Gas B exerts a pressure of 2.00 atm in a 2.00 dm^3 container. Gas C exerts a pressure of 3.00 atm in a 3.00 dm^3 container. What will be the total pressure if all three gases are placed in the 2.00 dm^3 container? The temperature is the same for all the gases.

OpenStudy (aaronq):

you can use \(P_1V_2=P_2V_2\) to find the P each gas will exert when compressed or expanded. Then add these up. for example, A : \((1\;atm)*(1L)=P_2(2L)\rightarrow P_2=\dfrac{(1\;atm)*(1L)}{(2L)}= 0.5\; atm\)

OpenStudy (jfraser):

and then since you need total pressure, add the individual pressures all together

OpenStudy (kira_yamato):

Thanks guys ^^